# On extension modules of finite homological dimension

### v1

@@ 2.8. Homological dimensions
The following class of $\mathcal{C}(R)$ are closed under short exact sequences.
  1. (1)
    ([10, Corollary 8.1.9]) The class of complexes in $\mathcal{C}_{(\sqsupset)}(R)$ with finite projective dimension.
  2. (2)
    ([10, Corollary 8.2.9])  The class of complexes in $\mathcal{C}_{(\sqsubset)}(R)$ with finite injective dimension.
  3. (3)
    ([6, Lemma 2.1.12 & Corollary 2.3.8]) The class of complexes in $\mathcal{C}_{(\sqsupset)}^{(\rm f)}(R)$ with finite Gorenstein dimension.
  4. (4)
    ([6, Lemma 3.2.12 & Corollary 6.2.5])  The class of complexes in $\mathcal{C}_{(\sqsubset)}(R)$ with finite Gorenstein injective dimension.
Theorem 2.2.
Let $R$ be a local ring and $\boldsymbol{X}$ be a complex.
  1. (1)
    ([10, Proposition 18.3.9]) If $\boldsymbol{X}\in\mathcal{C}_{(\sqsupset)}^{(\rm f)}(R)$, then $\operatorname{pd}_{R}(\boldsymbol{X})=\operatorname{pd}_{\widehat{R}}(\boldsymbol{X}\otimes_{R}\widehat{R}).$
  2. (2)
    ([10, Proposition 18.3.10]) If $\boldsymbol{X}\in\mathcal{C}_{(\sqsubset)}^{(\rm f)}(R)$, then $\operatorname{id}_{R}(\boldsymbol{X})=\operatorname{id}_{\widehat{R}}(\boldsymbol{X}\otimes_{R}\widehat{R}).$
  3. (3)
    ([23, 2.4]) If $\boldsymbol{X}\in\mathcal{C}_{(\square)}^{(\rm f)}(R)$, then $\operatorname{G-dim}_{R}(\boldsymbol{X})=\operatorname{G-dim}_{\widehat{R}}(\boldsymbol{X}\otimes_{R}\widehat{R}).$
  4. (4)
    ([10, Theorem 19.2.34]) If $\boldsymbol{X}\in\mathcal{C}_{(\square)}^{(\rm f)}(R)$ and $\operatorname{Gid}_{R}(\boldsymbol{X})<\infty$, then $\operatorname{Gid}_{R}(\boldsymbol{X})=\operatorname{Gid}_{\widehat{R}}(\boldsymbol{X}\otimes_{R}\widehat{R}).$
Proposition 2.3 ([4, Theorem 8.7]).
Let $R$ be a local ring and let $M$ be a finitely generated $R$-module. Then:
  1. (1)
    If $\operatorname{H\text{-}dim}_{R}(M)<\infty$, then $\operatorname{H\text{-}dim}_{R}(M)+0pt(M)=0pt(R)$.
  2. (2)
    For every integer $n\geq 0$, $\operatorname{H\text{-}dim}_{R}(\Omega^{n}(M))=\max\{\operatorname{H\text{-}dim}_{R}(M)-n,\,0\}.$
  3. (3)
    $\operatorname{G-dim}_{R}(M)\leq\operatorname{pd}_{R}(M)$ with equality whenever $\operatorname{pd}_{R}(M)$ is finite.
2.9. Semidualizing complexes.
An $R$-complex $\boldsymbol{C}\in\mathcal{C}_{(\square)}^{(\rm f)}(R)$ is semidualizing if the natural morphism $R\to\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{C},\boldsymbol{C})$ is an isomorphism in $\mathcal{D}(R)$. An $R$-module $C$ is said to be semidualizing if it corresponds to a semidualizing complex in $\mathcal{C}_{(\square)}^{(\rm f)}(R)$ ([7]).
2.10. Auslander categories
Let $\boldsymbol{C}$ be a semidualizing complex.
  1. (1)
    The Auslander class of $R$, denoted by $\mathcal{A}_{\boldsymbol{C}}(R)$, is the full subcategory of $\mathcal{D}(R)$ composed for all the complex $\boldsymbol{X}\in\mathcal{C}_{(\square)}(R)$ such that $\boldsymbol{C}\otimes_{R}^{\mathbf{L}}\boldsymbol{X}\in\mathcal{C}_{(\square)}(R)$ and the canonical map
    $\boldsymbol{X}\to\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{C},\boldsymbol{C}\otimes_{R}^{\mathbf{L}}\boldsymbol{X})$
    is an isomorphism in $\mathcal{D}(R)$.
  2. (2)
    The Bass class of $R$, denoted by $\mathcal{B}_{\boldsymbol{C}}(R)$, is the full subcategory of $\mathcal{C}_{(\square)}(R)$ composed for all the complexes $\boldsymbol{Y}\in\mathcal{C}_{(\square)}(R)$ such that $\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{C},\boldsymbol{Y})\in\mathcal{C}_{(\square)}(R)$ and the canonical map
    $\boldsymbol{C}\otimes_{R}^{\mathbf{L}}\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{C},\boldsymbol{Y})\to\boldsymbol{Y}$
    is an isomorphism in $\mathcal{D}(R)$.
Denote by $\mathcal{F}(R)$ and $\mathcal{I}(R)$, respectively, the subcategories of $\mathcal{D}(R)$ consisting of complexes isomorphic to homologically bounded complexes of flat $R$-modules and of injective $R$-modules. Then $\mathcal{F}(R)\subset\mathcal{A}_{\boldsymbol{C}}(R)$ and $\mathcal{I}(R)\subset\mathcal{B}_{\boldsymbol{C}}(R)$.
2.11. Dualizing complex
Assume that $R$ is a local ring. A dualizing complex $\boldsymbol{D}$ is a bounded complex of injective $R$-modules with finitely generated homologies such that the natural morphism $R\rightarrow\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{D},\boldsymbol{D})$ is an isomorphism in $\mathcal{D}(R)$.
The ring $R$ has a dualizing complex if and only if $R$ is a homomorphic image of a Gorenstein local ring ([6, A.8.3.2] and [24, Corollary 1.4]). In particular, if $R$ is complete, then $R$ admits a dualizing complex. If $\boldsymbol{D}$ and $\boldsymbol{D}^{\prime}$ are dualizing complexes, then $\boldsymbol{D}\simeq\Sigma^{n}\boldsymbol{D}$ for some integer $n$ ([6, A.8.3.3]).
If $\boldsymbol{X}$ is a complex in $\mathcal{C}_{(\square)}^{(\rm f)}(R)$, then $\boldsymbol{X}\simeq\operatorname{\mathbf{R}Hom}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{X},\boldsymbol{D}),\boldsymbol{D})$ ([6, A.8.5]).
Proposition 2.4.
Assume $R$ is a local ring and let $\boldsymbol{D}$ be a dualizing complex.
  1. (1)
    ([6, Corollary 3.1.11]) A complex $\boldsymbol{X}$ in $\mathcal{C}_{(\sqsupset)}^{(\rm f)}(R)$ has finite Gorenstein dimension if and only if $\boldsymbol{X}\in\mathcal{A}_{\boldsymbol{D}}(R)$.
  2. (2)
    ([9, Theorem 4.4]) A complex $\boldsymbol{X}$ in $\mathcal{C}_{(\sqsubset)}(R)$ has finite Gorenstein injective dimension if and only if $\boldsymbol{X}\in\mathcal{B}_{\boldsymbol{D}}(R)$.
A celebrated result that relates to the Auslander categories, and the subcategories $\mathcal{F}(R)$ and $\mathcal{I}(R)$ is the following, which is known as the Foxby Equivalence.
Theorem 2.5 ([6, Theorem 3.3.2]).
@@ 2.11. Dualizing complex
Assume that $R$ is a local ring with a dualizing complex $\boldsymbol{D}$. Let $\boldsymbol{X}$ and $\boldsymbol{Y}$ be complexes in $\mathcal{C}_{(\square)}(R)$. Then:
$\boldsymbol{X}\in\mathcal{A}_{\boldsymbol{D}}(R)\quad$ $\Longleftrightarrow\quad\boldsymbol{D}\otimes_{R}^{\mathbf{L}}\boldsymbol{X}\in\mathcal{B}_{\boldsymbol{D}}(R);$
$\boldsymbol{Y}\in\mathcal{B}_{\boldsymbol{D}}(R)\quad$ $\Longleftrightarrow\quad\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{D},\boldsymbol{Y})\in\mathcal{A}_{\boldsymbol{D}}(R);$
$\boldsymbol{X}\in\mathcal{F}(R)\quad$ $\Longleftrightarrow\quad\boldsymbol{D}\otimes_{R}^{\rm\mathbf{L}}\boldsymbol{X}\in\mathcal{I}(R);$
$\boldsymbol{Y}\in\mathcal{I}(R)\quad$ $\Longleftrightarrow\quad\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{D},\boldsymbol{Y})\in\mathcal{F}(R).$
In complementation to Theorem 2.2(4), one has the following.
Theorem 2.6 ([7, Proposition 5.9] & [10, Theorem 19.2.34]).
Let $R$ be a local ring with a dualizing complex and let $\boldsymbol{X}\in\mathcal{C}_{(\square)}^{(\rm f)}$. Then $\operatorname{Gid}_{R}(\boldsymbol{X})=\operatorname{Gid}_{\widehat{R}}(\boldsymbol{X}\otimes_{R}\widehat{R})$.
2.12. Stability results
We close this section with the following results, which will be used in the next section.
Theorem 2.7.
Assume $R$ is a local ring, and let $\boldsymbol{X}$ and $\boldsymbol{P}$ two complexes in $\mathcal{C}_{(\square)}^{(\rm f)}(R)$. Suppose that $\boldsymbol{P}$ is non-acylic and that $\operatorname{pd}_{R}(\boldsymbol{P})<\infty$. Then:
  1. (1)
    ([23, Theorems 5.6 & 5.8]) $\operatorname{H-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{P},\boldsymbol{X}))<\infty\Longleftrightarrow\operatorname{H-dim}_{R}(\boldsymbol{X})<\infty$.
  2. (2)
    ([23, Theorems 5.1 & 5.7]) $\operatorname{H-dim}_{R}(\boldsymbol{P}\otimes_{R}^{\mathbf{L}}\boldsymbol{X})<\infty$Longleftrightarrow $\operatorname{H-dim}_{R}(\boldsymbol{X})<\infty$.
Theorem 2.8.
([9, Corollary 6.4]) Assume $R$ is a local ring with a dualizing complex $\boldsymbol{D}$. Then for a complex $\boldsymbol{X}\in\mathcal{C}_{(\square)}^{(\rm f)}(R)$ one has:
$\operatorname{Gid}_{R}(\boldsymbol{X})<\infty\Longleftrightarrow\operatorname{G-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{X},\boldsymbol{D}))<\infty.$
Theorem 2.9.
Assume $R$ is a local ring, and let $\boldsymbol{Y}$ and $\boldsymbol{P}$ be complexes in $\mathcal{C}_{(\square)}^{(\rm f)}(R)$. Suppose that $\boldsymbol{P}$ is non-acyclic and that $\operatorname{pd}_{R}(\boldsymbol{P})<\infty$. In the case $\operatorname{H-dim}_{R}=\operatorname{G-dim}_{R}$, assume further that $R$ has a dualizing complex.
  1. (1)
    If $\operatorname{Hid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{P},\boldsymbol{Y}))<\infty$, then $\operatorname{Hid}_{R}(\boldsymbol{Y})<\infty$.
  2. (2)
    If $\operatorname{Hid}_{R}(\boldsymbol{Y}\otimes_{R}^{\mathbf{L}}\boldsymbol{P})<\infty$, then $\operatorname{Hid}_{R}(\boldsymbol{Y})<\infty$.
Proof.
If $\operatorname{H-dim}=\operatorname{pd}$, it follows from [8, Lemmas 6.2.10 & 6.2.15]. On the other hand, if $\operatorname{H-dim}=\operatorname{G-dim}$, then the theorem follows from [32, Proposition 4.1(b)] and Proposition 2.4(2). ∎
3. Finite homological dimensions of Ext of complexes
In this section, we derive results regarding the finiteness of homological dimension of Ext of complexes, and conclude the corresponding results for Ext of modules. The motivation arises from the [37, Section 6], where the respective authors obtained a result concerning finiteness of the injective dimension of Ext of complexes ([37, Proposition 4.2]) using [37, Lemma 4.1] and Bass Series, and subsequently derived the corresponding result for Ext of modules. Inspired by their approach, we first derive counteparts of the lemma mentioned considering other homological dimensions and then use Poincaré Series or stability results for concluding our results.
Lemma 3.1.
Let
$\boldsymbol{X}:\cdots\longrightarrow X_{l+1}\xrightarrow{\partial_{l+1}}X_{l}\xrightarrow{\partial_{l}}X_{l-1}\rightarrow\cdots$
be a homologically bounded complex. If $\operatorname{Gid}_{R}(\operatorname{H}_{i}(\boldsymbol{X}))<\infty$ for all $i,$ then $\operatorname{Gid}_{R}(\boldsymbol{X})<\infty$.
Proof.
The proof is the same as the proof of [37, Lemma 4.1] but now replacing $\operatorname{Gid}$ by $\operatorname{id}$, considering the complex $\boldsymbol{I}$ of the proof as a resolution of Gorenstein injective modules of $\operatorname{H}_{0}(\boldsymbol{X})$, and using Theorem 2.1(4). ∎
Lemma 3.2.
Let
$\boldsymbol{X}:\cdots\longrightarrow X_{l+1}\xrightarrow{\partial_{l+1}}X_{l}\xrightarrow{\partial_{l}}X_{l-1}\rightarrow\cdots$
be a complex in $\mathcal{C}_{(\square)}(R)$. Suppose further that $\boldsymbol{X}\in\mathcal{C}^{(\rm{f})}(R)$ in the case $\operatorname{H-dim}_{R}=\operatorname{G-dim}$. If $\operatorname{H-dim}_{R}(\operatorname{H}_{i}(\boldsymbol{X}))<\infty$ for all $i,$ then $\operatorname{H-dim}_{R}(\boldsymbol{X})<\infty$.
Proof.
Let $n$ be the number of non-zero integers $i$ such that $\operatorname{H}_{i}(\boldsymbol{X})\not=0$. If $n=0,$ then $\boldsymbol{X}$ is acyclic and hence $\operatorname{H-dim}_{R}(\boldsymbol{X})=-\infty<\infty$. Now, we show by induction that the lemma holds when $n\geq 1$.
If $n=1,$ then there exists an unique integer $i$ such that $\operatorname{H}_{i}(\boldsymbol{X})\not=0$. Without loss of generality, we may assume $i=0$. Set $Z=\operatorname{Coker}(\partial_{1})$. Let
$\boldsymbol{Y}:\cdots\rightarrow 0\rightarrow Z\xrightarrow{\tilde{\partial}_{0}}X_{-1}\xrightarrow{\partial_{-1}}X_{-2}\rightarrow\cdots,$
where ~$\tilde{\partial}_{0}$ is the map induced of $\partial_{0}$. The natural homomorphism $\boldsymbol{X}\to\boldsymbol{Y}$ is a quasi-isomorphism, so $\boldsymbol{X}\simeq\boldsymbol{Y}$ in $\mathcal{D}(R)$. Also, we may note that the natural homomorphism $\operatorname{H}_{0}(\boldsymbol{X})\to\boldsymbol{Y}$, is a quasi-isomorphism, so $\operatorname{H}_{0}(\boldsymbol{X})\simeq\boldsymbol{Y}$ in $\mathcal{D}(R)$. Thus, $\operatorname{H}_{0}(\boldsymbol{X})\simeq\boldsymbol{X}$ in $\mathcal{D}(R)$. As $\operatorname{H-dim}_{R}(\operatorname{H}_{0}(\boldsymbol{X}))<\infty,$ it follows that $\operatorname{H-dim}_{R}(\boldsymbol{X})<\infty$.
Suppose $n>1$. We may suppose that $\sup\{i\in\mathbb{Z}:\operatorname{H}_{i}(\boldsymbol{X})\neq 0\}=0$. As above, if $Z=\operatorname{Coker}(\partial_{1})$ and $\boldsymbol{Y}$ is the complex
$\boldsymbol{Y}:\cdots\rightarrow 0\rightarrow Z\xrightarrow{\tilde{\partial}_{0}}X_{-1}\xrightarrow{\partial_{-1}}X_{-2}\rightarrow\cdots,$
then $\boldsymbol{X}\simeq\boldsymbol{Y}$ in $\mathcal{D}(R)$. Since $\operatorname{H-dim}_{R}(\operatorname{H}_{0}(\boldsymbol{X}))<\infty$, then there exists a bounded complex
$\boldsymbol{T}:\cdots\rightarrow T_{i+1}\rightarrow T_{i}\rightarrow\cdots\rightarrow T_{1}\rightarrow T_{0}\rightarrow 0$
such that $\operatorname{H-dim}_{R}(T_{i})=0$ for all $i\geq 0$, and
$\operatorname{H}_{i}(\boldsymbol{T})=\begin{cases}\operatorname{H}_{0}(\boldsymbol{X}),&\text{if }i=0\\ 0,&\text{if }i>0.\end{cases}$
Let $\boldsymbol{C}$ be the complex
$\boldsymbol{C}:\cdots\rightarrow T_{1}\rightarrow T_{0}\rightarrow Z\xrightarrow{\tilde{\partial}_{0}}X_{-1}\xrightarrow{\partial_{-1}}X_{-2}\rightarrow\cdots,$
where $T_{0}\to Z$ is the composition of the maps $T_{0}\to\operatorname{H}_{0}(X)$ and $\operatorname{H}_{0}(X)\to Z$. Let us consider that the complexes $\boldsymbol{Y}$ and $\boldsymbol{C}$ are indexing such that $Y_{-1}=Z$ and $C_{0}=T_{0}$. Then we have naturally an exact sequence
$0\to\boldsymbol{Y}\to\boldsymbol{C}\to\boldsymbol{T}\to 0$
Note that $\operatorname{H}_{i}(\boldsymbol{C})=0$ for all $i\geq-1$, and $\operatorname{H}_{i}(\boldsymbol{C})=\operatorname{H}_{i+1}(\boldsymbol{X})$ for $i\leq-2$. Hence, the number of non-zero homologies of $\boldsymbol{C}$ is $n-1$. Thus, by induction, $\operatorname{H-dim}_{R}(\boldsymbol{C})<\infty$. Then from Theorem 2.1(1) or (3), we see that $\operatorname{H-dim}_{R}(\boldsymbol{Y})<\infty$. Thus, since $\boldsymbol{X}\simeq\boldsymbol{Y}$, we see that $\operatorname{H-dim}_{R}(\boldsymbol{X})<\infty$. ∎
Theorem 3.3.
Assume $R$ is local, and let $\boldsymbol{M}$ and $\boldsymbol{N}$ be non-acyclic complexes in $\mathcal{C}_{(\square)}^{\rm(f)}(R)$ such that $\operatorname{pd}_{R}(\boldsymbol{M})<\infty$.
@@ 3. Finite homological dimensions of Ext of complexes
We know that $\operatorname{Ext}_{R}^{i}(\boldsymbol{M},\boldsymbol{N})=\operatorname{H}_{-i}(\mathbf{R}\mathrm{Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))$ for all $i$. From [6, A.4.4 and A.5.4], ones sees that $\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N})\mathcal{C}_{(\square)}^{\rm(f)}(R)$. By Lemma 3.2 (resp. Lemma 3.1), we conclude $\operatorname{H-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$ (resp. $\operatorname{Gid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$). It follows from Theorem 2.7(1) (resp. Theorem 2.9(1)) that $\operatorname{H-dim}_{R}(\boldsymbol{N})<\infty$. (resp. $\operatorname{Gid}_{R}(\boldsymbol{N})<\infty$), concluding (1) (resp. (2)). ∎
@@
Item (2) of the next theorem provides a positive answer to the Gorenstein version of [18, Question 4.5].
@@
Assume $R$ is local, and let $\boldsymbol{M}$ and $\boldsymbol{N}$ be non-acyclic complexes in $\mathcal{C}_{(\square)}^{(\rm f)}(R)$. Suppose that $\operatorname{id}_{R}(\boldsymbol{N})<\infty$. In the case $\operatorname{H-dim}=\operatorname{G-dim}$, assume further that $R$ has a dualizing complex.
@@
We know that $\operatorname{Ext}_{R}^{i}(\boldsymbol{M},\boldsymbol{N})=\operatorname{H}_{-i}(\mathbf{R}\mathrm{Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))$. Additionally, by [6, A.4.4 and A.5.2], we see that $\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N})\in\mathcal{C}_{(\square)}^{(\rm f)}(R)$. In the case (1) (resp. (2)), by assumption, $\operatorname{H-dim}_{R}(\operatorname{H}_{i}(\mathbf{R}\mathrm{Hom}_{R}(\boldsymbol{M},\boldsymbol{N})))<\infty$ (resp. $\operatorname{Gid}_{R}(\operatorname{H}_{i}(\mathbf{R}\mathrm{Hom}_{R}(\boldsymbol{M},\boldsymbol{N})))<\infty$) for all $i$. Thus, by Lemma 3.1 (resp. Lemma 3.2), we have that $\operatorname{H-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$ (resp. $\operatorname{Hid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$) for all $i$.
The conclusion of item (1) for $\operatorname{H-dim}_{R}=\operatorname{pd}_{R}$ it follows from [8, Lemma 6.2.12].
It is remaining to prove (1) for $\operatorname{H-dim}_{R}=\operatorname{G-dim},$ and (2). For proving both cases, either by Theorems 2.2, or by assumption and Theorem 2.6, we can let $\boldsymbol{D}$ be a dualizing complex of $R$. Since $\operatorname{id}_{R}(\boldsymbol{N})<\infty$, then $\boldsymbol{N}\in\mathcal{B}_{\boldsymbol{D}}(R)$ and hence $\operatorname{pd}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{D},\boldsymbol{N}))<\infty$ by the Foxby equivalence.
@@
$\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N})$ $\simeq\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{D}\otimes_{R}^{\mathbf{L}}\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{D},\boldsymbol{N}))$ because $\text{because }\boldsymbol{N}\in\mathcal{B}_{\boldsymbol{D}}(R)$
@@
In the case (1) for $\operatorname{H-dim}_{R}=\operatorname{G-dim}_{R}$, we have that $\operatorname{G-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$, concluding from Theorem 2.7(2) that $\operatorname{G-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{D}))<\infty$. Hence $\operatorname{Gid}_{R}(\boldsymbol{M})<\infty$ by Theorem 2.8.
As to (2), we have that $\operatorname{Gid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$. Then from the last isomorphism and Theorem 2.9(2), we conclude that $\operatorname{Gid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{D}))<\infty$. Since $M\simeq\operatorname{\mathbf{R}Hom}_{R}(\operatorname{\mathbf{R}Hom}_{R}(M,\boldsymbol{D}),\boldsymbol{D})$ by [6, A.8.5], then it follows from Theorem 2.8 that $\operatorname{G-dim}_{R}(\boldsymbol{M})<\infty$.
@@
When $\operatorname{H-dim}_{R}=\operatorname{pd}_{R}$, note that Corollary 3.7(1) recovers the celebrated theorem given by Foxby in [15, Corollary 4.4], which states that if there exists a (finitely generated) non-zero $R$-module $M$ with both finite injective dimension and projective dimension, then $R$ is Gorenstein. On the other hand, via Corollary 3.7(2), we obtain the following Gorenstein criterion that generalizes [37, Corollary 4.4] and [18, Corollary 4.7].

### v2

@@ 2. Finite homological dimensions of Ext of complexes
Assume $R$ is local, and let $\boldsymbol{M}$ and $\boldsymbol{N}$ be non-acyclic complexes in $\mathcal{D}_{\square}^{\rm f}(R)$ such that $\operatorname{pd}_{R}(\boldsymbol{M})<\infty$.
@@
We know that $\operatorname{Ext}_{R}^{i}(\boldsymbol{M},\boldsymbol{N})=\operatorname{H}_{-i}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))$ for all $i$. From [6, A.4.4 and A.5.4], one sees that $\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N})\in\mathcal{D}_{\square}^{\rm f}(R)$. By Lemma 2.1(2) (resp. Lemma 2.1(1)), we conclude $\operatorname{H-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$ (resp. $\operatorname{Gid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$). It follows from [19, Theorems 5.6 & 5.8] (resp. [28, Proposition 4.1(b)] and [9, Theorem 4.4]) that $\operatorname{H-dim}_{R}(\boldsymbol{N})<\infty$ (resp. $\operatorname{Gid}_{R}(\boldsymbol{N})<\infty$), concluding (1) (resp. (2)). ∎
@@
Item (2) of the next theorem provides a positive answer to the Gorenstein version of [15, Question 4.5].
@@
Assume $R$ is local, and let $\boldsymbol{M}$ and $\boldsymbol{N}$ be non-acyclic complexes in $\mathcal{D}_{\square}^{\rm f}(R)$. Suppose that $\operatorname{id}_{R}(\boldsymbol{N})<\infty$. In the case $\operatorname{H-dim}=\operatorname{G-dim}$, assume further that $R$ has a dualizing complex.
@@
We know that $\operatorname{Ext}_{R}^{i}(\boldsymbol{M},\boldsymbol{N})=\operatorname{H}_{-i}(\mathbf{R}\mathrm{Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))$. Additionally, by [6, A.4.4 and A.5.2], we see that $\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N})\in\mathcal{D}_{\square}^{\rm f}(R)$. In the case (1), for all $i$, $\operatorname{H-dim}_{R}(\operatorname{H}_{i}(\mathbf{R}\mathrm{Hom}_{R}(\boldsymbol{M},\boldsymbol{N})))<\infty$. Thus, by Lemma 2.1(1), we have that $\operatorname{H-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$. As to (2), one considers Lemma 2.1(2) instead so that $\operatorname{Gid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$.
The conclusion of item (1) for $\operatorname{H-dim}_{R}=\operatorname{pd}_{R}$ follows from [8, Lemma 6.2.12].
It remains to prove (1) for $\operatorname{H-dim}_{R}=\operatorname{G-dim},$ and (2). For proving both, either by [19, 2.4], or by assumption, [7, Proposition 5.9] and [10, Theorem 19.2.34], we can let $\boldsymbol{D}$ be a dualizing complex of $R$. Since $\operatorname{id}_{R}(\boldsymbol{N})<\infty$, then $\boldsymbol{N}\in\mathcal{B}(R)$ and hence $\operatorname{pd}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{D},\boldsymbol{N}))<\infty$ by the Foxby equivalence.
@@
$\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N})$ $\simeq\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{D}\otimes_{R}^{\mathbf{L}}\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{D},\boldsymbol{N}))$ $\text{because }\boldsymbol{N}\in\mathcal{B}(R)$
@@
In the case (1) for $\operatorname{H-dim}_{R}=\operatorname{G-dim}_{R}$, we have that $\operatorname{G-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$, concluding from [19, Theorems 5.1] that $\operatorname{G-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{D}))<\infty$. Hence $\operatorname{Gid}_{R}(\boldsymbol{M})<\infty$ by [9, Corollary 6.4].
As to (2), we have that $\operatorname{Gid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$. It follows from the last isomorphism, [28, Proposition 4.1(b)] and [9, Theorem 4.4] that $\operatorname{Gid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{D}))<\infty$. Since $M\simeq\operatorname{\mathbf{R}Hom}_{R}(\operatorname{\mathbf{R}Hom}_{R}(M,\boldsymbol{D}),\boldsymbol{D})$ by [6, A.8.5], then it follows from [9, Corollary 6.4] that $\operatorname{G-dim}_{R}(\boldsymbol{M})<\infty$. ∎
@@
When $\operatorname{H-dim}_{R}=\operatorname{pd}_{R}$, note that Corollary 2.6(1) recovers the celebrated theorem of Foxby in [14, Corollary 4.4], which states that if there exists a (finitely generated) non-zero $R$-module $M$ with both finite injective dimension and projective dimension, then $R$ is Gorenstein. On the other hand, via Corollary 2.6(2), we obtain the following Gorenstein criterion that generalizes [15, Corollary 4.7] and [32, Corollary 4.4].